An object, that remains at rest, is acted upon by three forces F1, F2 and F3. The component forms of F1 and F2 are given by

F1 =ai + bj and F2 = ci + dj

where a = -116, b = 50, c = 129 and d = -195. If the component form of F3 is of the form

F3 = xi + yj

then what is the value of x?
1
Expert's answer
2011-10-06T12:42:23-0400
As objectremains at rest then
Fx=Fy=0
Then a i+b i+x i =0=-116 i+129 i+x i =13 i+x i =0 => x=-13
please note that i - vector